The sum of the digit of a two digit number is 9. the ten digit number is 3 more than twice the one digit.
Answer by mananth(16949)

let the number be xy
x+y =9
x=2y+3
plug value of x
in first equation
2y+3+y = 9
3y+3=9
3y=9-3
3y=6
y=2
...
x+y =9
y =2
so x = 7
number = 72
...
m.ananth@hotmail.ca